You have to design a logical circuit for a 4-bit composite number using K-MAP (SOP minimization). Provide the original and simplified Boolean expressions in SOP form. Covert SOP to POS using the K-Map.


1
Expert's answer
2021-05-05T07:41:29-0400

Logic circuit for 4 bit composite number is

"Y=AC+ABC + \bar{A}BC+AB+D"

Simplifying the given boolean function

"Y=AC+ABC + \bar{A}BC+AB+D"

"Y=AC+BC +(A \bar{A})+AB+D"


SOP "\implies Y=AC+BC + AB+D"


POS "\implies Y=(\bar{A}+ \bar{C})(\bar{B}+ \bar{C})(\bar{A}+ \bar{B})(\bar{D})"


Implementation of SOP in a logic diagram


"Y=AC+BC + AB+D"